A magnet fiel is at 30 degrees to the axis of a coil with 300 turns and radius of 4cm. If the magnetic field changes at 85 Tesla per seconds, find the induced EMF. If teh resistance of the coil is 200 olms, find the induced current in the coil. show your work
Answer #1

The cross sectional area of the coil is,

A=pi(r^2)

=pi(0.04 m)^2

=0.005024 m^2

The induced emf is,

E = NAcos(thita)(dB/dt)

=(300)(0.005024)(cos30)(85)

=110.9 V

The current induced is,

I= V/R

=(110.9 V)/(200 ohm)

=0.554 A